Barbell Company reported net operating income (NOI) equal to $180,000 this year. Examination of the company's balance sheet and income statement shows that the tax rate was 35 percent, the depreciation expense was $50,000, $140,000 was invested in assets during the year, and invested capital currently is $1,200,000. If Barbell' average after-tax cost of funds is 9 percent, what is the firm's EVA?
Answer #1
EVA = NOPAT-(Invested capital*WACC)
NOPAT = NOI*(1-t) = 180000*(1-35%) = $      1,17,000
Invested capital = 1200000-140000 = $    10,60,000
EVA = 117000-1060000*9% = $ 21,600